# Optimal Algorithm for Multiple-Processor Multitask Scheduling

### v1

@@ Abstract
The efficient scheduling of multi-task jobs across multiprocessor systems has become increasingly critical with the rapid expansion of computational systems. This challenge, known as Multiprocessor Multitask Scheduling (MPMS), is essential for optimizing the performance and scalability of applications in fields such as cloud computing and deep learning. In this paper, we study the MPMS problem under both deterministic and stochastic models, where each job is composed of multiple tasks and can only be completed when all its tasks are finished. We introduce $\mathsf{NP}$-$\mathsf{SRPT}$, a non-preemptive variant of the Shortest Remaining Processing Time (SRPT) algorithm, designed to accommodate scenarios with non-preemptive tasks. Our algorithm achieves a competitive ratio of $\ln\alpha+\beta+1$ for minimizing response time, where $\alpha$ represents the ratio of the largest to the smallest job workload, and $\beta$ captures the ratio of the largest non-preemptive task workload to the smallest job workload. We further establish that this competitive ratio is order-optimal when the number of processors is fixed. For stochastic systems modeled as M/G/N queues, where job arrivals follow a Poisson process and task workloads are drawn from a general distribution, we prove that $\mathsf{NP}$-$\mathsf{SRPT}$ achieves asymptotically optimal mean response time as the traffic intensity $\rho$ approaches $1$, assuming the task size distribution has finite support. Moreover, the asymptotic optimality extends to cases with infinite task size distributions under mild probabilistic assumptions, including the standard M/M/N model. Experimental results validate the effectiveness of $\mathsf{NP}$-$\mathsf{SRPT}$, demonstrating its asymptotic optimality in both theoretical and practical settings.
1 Introduction
@@
Optimizing the response time of single-task jobs has been considered both in offline and online scenarios. If preemption is allowed, the Shortest Remaining Processing Time (SRPT) discipline is shown to be optimal in single machine environment. Many generalizations of this basic formulation become NP-hard, for example, minimizing the total response time in non-preemptive single machine model and preemptive model with two machines [1]. When jobs arrive online, no information about jobs is known to the algorithm in advance, several algorithms with logarithmic competitive ratios are proposed in various settings [6, 1]. On the other hand, while SRPT minimizes the mean response time sample-path wise, it requires the knowledge of remaining job service time. Gittins proved that the Gittins index policy minimizes the mean response time in an M/G/1 queue, which only requires the access to the information about job size distribution [7].
@@
Multi-processor Multitask Scheduling (MPMS) [9], where a job is considered to be completed only when all the tasks within the job are finished, addresses the allocation of jobs and tasks, each with potentially varying computational demands, to a set of processors in a manner that optimizes overall system performance metrics. Though much progresses have been made in single-task job scheduling, there is a lack of theoretical understanding regarding MPMS. The extension of traditional scheduling principles to multi-task systems, presents significant challenges. The heterogeneity of task durations and the dynamic nature of job arrivals seems further complicate the analysis. Thus a natural question that arises is,
@@
  • In Section 3, we propose $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm [13], for minimizing the total response time. $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm achieves a competitive ratio of $\ln\alpha+\beta+1$, where $\alpha$ is the maximum-to-minimum job workload ratio, $\beta$ represents the ratio between maximum non-preemptive task workload and minimum job workload. It can be shown that no $o(\ln\alpha+\beta)$-competitive algorithm exists when the number of machines is constant. In addition, $O(\ln\alpha+\beta^{1-\varepsilon})$ is the best possible competitive ratio for the class of work-conserving algorithms.
@@
  • Besides the worst case relative ratio above, we further prove our main result, absolute performance guarantees for $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm under certain probabilistic structure on the input instances, in which the relevant and remaining workload bound established for the adversarial inputs contributes significantly to the stochastic analysis. Assuming that jobs arrive according to a Poisson process, i.e., in M/G/N system, in Section 4 we prove that the average response time incurred by $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm is asymptotic optimal when load $\rho\rightarrow 1$, as long as the task size distribution has finite support. As shown in Section 5, the assumption of finite task workload can be removed for exponentially distributed job size, i.e., M/M/N, together with other job size distributions with certain properties on the tail of the distribution, by establishing an upper bound on the busy period in M/G/N.
@@
In addition, we also further validate the optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ through experiments in Section 6, confirming the efficacy of the algorithm.
1.2 Related Work
Single and multiple task scheduling.
@@ 1.2 Related Work
There has been a large literature on single-task job scheduling, with parallel developments taking place in competitive analysis and queuing theory. For example, recently Hong and Scully [14] showed Gittins’s heavy-traffic optimality in G/G/N. However, little is known about multitask scheduling. Scully et. al [15] presented the first theoretical analysis of single-processor multitask scheduling problem, and gave an optimal policy that is easy to compute for batch arrival, together with the assumption that the processing time of tasks satisfies the aged Pareto distributions. Sun et al. [16] studied the multitask scheduling problem when all the tasks are of unit size, and proved that among causal and non-preemptive policies, fewest unassigned tasks first (FUT) policy, earliest due date first (EDD) policy, and first come first serve (FCFS) are near delay-optimal in distribution (stochastic ordering) for minimizing the metric of average delay, maximum lateness and maximum delay respectively. Wang et. al [3] established results of asymptotic independence of queues in multitask multi-processor system by developing a new technique named Poisson oversampling. To model the scenario when the scheduler has incomplete information about the job size, Scully et. al [17] introduced the multistage job model and proposed an optimal scheduling algorithm for multistage job scheduling in M/G/1 queue. The closed-form expression of the mean response time is also given for the optimal scheduler. In addition to the aforementioned work, there are also studies that further extend the understanding of scheduling by considering parallelizable jobs represented as Directed Acyclic Graphs (DAGs) [18].
Performance and optimality of SRPT and its variants.
@@
While SRPT is optimal at minimizing average response time in single-server systems [19], its performance is suboptimal in multi-server environments. However, SRPT remains a highly regarded method in the context of multi-server systems. It has shown that the competitive ratio of SRPT is in logarithmic order and is the best possible [1]. With the stochastic arrivals in M/G/k systems, SRPT is shown to be optimal in heavy traffic scenarios [20]. Another notable contribution by [21] introduces the Intermediate-SRPT algorithm, which addresses jobs with intermediate parallelizability and establishes an optimal $O(\log P)$-competitive ratio concerning average flow time, where $P$ represents the ratio of maximum to minimum job sizes.
Paper organization.
@@
The remainder of this paper is organized as following. We introduce the problem definition, notations and necessary background in Section 2. In Section 3 we formally present $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, together with the analysis of its competitive ratio and lower bounds. Section 4 is devoted to the proof of the asymptotic optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ in heavy traffic regime, and the optimality is extended to infinite task size in Section 5. We conduct the experimental validation in Section 6, and conclude our work in Section 7.
2 Model and preliminaries
Deterministic Model.
@@ 2 Model and preliminaries
[…]
@@
Stochastic Model.
In the stochastic setting, we assume that jobs arrive into the system according to a Poisson process with rate $\lambda$. Job processing times are i.i.d distributed with probability density function $f(\cdot)$. Formally, we consider a sequence of M/G/N queues indexed by $n$, the traffic intensity of the $n$-th system is equal to $\rho^{(n)}=\lambda^{(n)}\cdot\mathbbm{E}[p^{(n)}_{i}]$, where $\lambda^{(n)}$ denotes the arrival rate of the $n$-th Poisson arrival process, job workload distribution has a density function of $f^{(n)}(\cdot)$. Stability of the queuing systems requires that $\rho^{(n)}<1$ for $\forall n$. As standard in the literature, we assume that $\rho^{(n)}\rightarrow 1$ when $n\rightarrow\infty$. In this paper, we further assume that the probability density function $f^{(n)}(\cdot)$ is continuous. For notational convenience, we will suppress index $n$ whenever it is clear from the context.
The stochastic analysis in this paper relies heavily on the concept of busy period in M/G/1, which is defined as following.
Definition 3 (Busy Period in M/G/1 [22]).
Busy period in M/G/1 is defined to be the longest time interval in which no machines are idle.
We use $\mathsf{B}(w)$ to denote the expected length of a busy period in M/G/1, started by a workload of $w$. It can be seen that $\mathsf{B}(\cdot)$ is an additive function [22], i.e., $\mathsf{B}(w_{1}+w_{2})=\mathsf{B}(w_{1})+\mathsf{B}(w_{2})$ for any independent initial workload $w_{1},w_{2}$, since a busy period in M/G/1 with initial workload of $w_{1}+w_{2}$ can be regarded as a busy period started by initial workload $w_{2}$, following a busy period started by initial workload $w_{1}$. Moreover, for M/G/1 queue, the length of a busy period with initial workload of $w$ and load $\rho$ is shown to be equal to $\mathsf{B}(w)=\mathbbm{E}[w]/(1-\rho)$ [22].
SOAP policy.
@@
A SOAP policy is monotonic if its rank function is monotonic in age $a$. In this paper, a tie-breaking rule, First-Come First-Served (FCFS) is applied to determine the order of service in the event of a tie.
3 $\mathsf{NP}$-$\mathsf{SRPT}$ Algorithm and Competitive Ratio Analysis

### v2

# On Resolving Non-Preemptivity in Multitask Scheduling: An Optimal Algorithm in Deterministic and Stochastic Worlds

@@ Abstract
The efficient scheduling of multi-task jobs across multiprocessor systems has become increasingly critical with the rapid expansion of computational systems. This challenge, known as Multiprocessor Multitask Scheduling (MPMS), is essential for optimizing the performance and scalability of applications in fields such as cloud computing and deep learning. In this paper, we study the MPMS problem under both deterministic and stochastic models, where each job is composed of multiple tasks and can only be completed when all its tasks are finished. We introduce $\mathsf{NP}$-$\mathsf{SRPT}$, a non-preemptive variant of the Shortest Remaining Processing Time (SRPT) algorithm, designed to accommodate scenarios with non-preemptive tasks. Our algorithm achieves a competitive ratio of $\ln\alpha+\beta+1$ for minimizing response time, where $\alpha$ represents the ratio of the largest to the smallest job workload, and $\beta$ captures the ratio of the largest non-preemptive task workload to the smallest job workload. We further establish that this competitive ratio is order-optimal when the number of processors is fixed. For stochastic systems modeled as $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ queues, where job arrivals follow a Poisson process and task workloads are drawn from a general distribution, we prove that $\mathsf{NP}$-$\mathsf{SRPT}$ achieves asymptotically optimal mean response time as the traffic intensity $\rho$ approaches $1$, assuming the task size distribution has finite support. Moreover, the asymptotic optimality extends to cases with infinite task size distributions under mild probabilistic assumptions, including the standard $\mathsf{M}$$\mathsf{M}$$\mathsf{N}$ model. Finally, we extend the analysis to the practical setting of unknown job sizes, proving that non-preemptive adaptations of the $\mathsf{M\text{-}Gittins}$ and $\mathsf{M\text{-}SERPT}$ policies achieve asymptotic optimality and near-optimality, respectively, for a broad class of job size distributions. Experimental results validate the effectiveness of $\mathsf{NP}$-$\mathsf{SRPT}$, demonstrating its asymptotic optimality in both theoretical and practical settings.
I Introduction
@@
Optimizing the response time of single-task jobs has been considered both in offline and online scenarios. If preemption is allowed, the Shortest Remaining Processing Time (SRPT) discipline is shown to be optimal in single machine environment. Many generalizations of this basic formulation become NP-hard, for example, minimizing the total response time in non-preemptive single machine model and preemptive model with two machines [23]. When jobs arrive online, no information about jobs is known to the algorithm in advance, several algorithms with logarithmic competitive ratios are proposed in various settings [3, 23]. On the other hand, while SRPT minimizes the mean response time sample-path wise, it requires the knowledge of remaining job service time. Gittins proved that the Gittins index policy minimizes the mean response time in an $\mathsf{M}$$\mathsf{G}$/11 queue, which only requires the access to the information about job size distribution [11].
Figure 1: Illustration of the multi-task multi-processor scheduling model. Jobs consist of multiple (possibly non-preemptive) tasks, which are placed into a single centralized global queue managed by the dispatcher. Servers fetch tasks from the global queue for execution, and a job is considered complete only when all its constituent tasks have finished.
@@
Multi-processor Multitask Scheduling (MPMS) [42], as shown in Figure 1, where a job is considered to be completed only when all the tasks within the job are finished, addresses the allocation of jobs and tasks, each with potentially varying computational demands, to a set of processors in a manner that optimizes overall system performance metrics. Though much progresses have been made in single-task job scheduling, there is a lack of theoretical understanding regarding MPMS. The extension of traditional scheduling principles to multi-task systems, presents significant challenges. The heterogeneity of task durations and the dynamic nature of job arrivals seems further complicate the analysis. Thus a natural question that arises is,
@@
  • In Section III, we propose $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, for minimizing the total response time. $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm achieves a competitive ratio of $\ln\alpha+\beta+1$, where $\alpha$ is the maximum-to-minimum job workload ratio, $\beta$ represents the ratio between maximum non-preemptive task workload and minimum job workload. It can be shown that no $o(\ln\alpha+\beta)$-competitive algorithm exists when the number of machines is constant. In addition, $O(\ln\alpha+\beta^{1-\varepsilon})$ is the best possible competitive ratio for the class of work-conserving algorithms.
@@
  • Besides the worst case relative ratio above, we further prove our main result, absolute performance guarantees for $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm under certain probabilistic structure on the input instances, in which the relevant and remaining workload bound established for the adversarial inputs contributes significantly to the stochastic analysis. Assuming that jobs arrive according to a Poisson process, i.e., in $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ system, in Section IV we prove that the average response time incurred by $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm is asymptotic optimal when load $\rho\rightarrow 1$, as long as the task size distribution has finite support. As shown in Section V, the assumption of finite task workload can be removed for exponentially distributed job size, i.e., $\mathsf{M}$$\mathsf{M}$$\mathsf{N}$, together with other job size distributions with certain properties on the tail of the distribution, by establishing an upper bound on the busy period in $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$.
  • In Section VI, we address the practical challenge of scheduling jobs with unknown sizes. The analysis is extended to show that non-preemptive versions of policies designed for this setting, namely $\mathsf{M\text{-}Gittins}$ and $\mathsf{M\text{-}SERPT}$, are asymptotically optimal and 2-competitive, respectively, in the heavy-traffic limit for a broad class of job size distributions.
@@
In addition, we also further validate the optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ through experiments in Section VII, confirming the efficacy of the algorithm.
I-B Analytical challenges in MPMS and our approach
In many respects, preemptive model are significantly more tractable due to the inherent decomposability of the model. In preemptive-resume disciplines, the highest-priority class can be modeled as a standard $\mathsf{M}/\mathsf{G}/1$ queue, while lower classes can be represented as $\mathsf{M}/\mathsf{G}/1$ queues with busy periods generated by higher-priority classes. This allows the well-developed theory of $\mathsf{M}/\mathsf{G}/1$ queues to be applied iteratively, requiring only aggregate statistics of higher-priority workloads [13]. Such tractability sharply contrasts with non-preemptive scheduling, including the multitask problem considered in this paper, where blocking implies that high-priority jobs may be delayed by the completion of lower-priority non-preemptive tasks already in service, thereby inducing a tight coupling between all priority levels. Thus the performance of a certain class arrival depends not only on the higher-priority classes but also on the detailed instantaneous states of lower-priority jobs, including their remaining service times. This dependence eliminates the possibility of recursive analysis and necessitates richer state descriptions and more complex techniques, such as multidimensional Markov chains [7].
In multi-processor environments with a finite number of job classes, the analysis quickly becomes intractable due to state-space explosion. To fully describe the system, one must specify the class of the job occupying each server. Even with only five servers and five priority classes, the number of possible states already exceeds one billion [7]. In multi-task systems, this complexity intensifies, as the system state must additionally capture the internal progress of every active job. The resulting multiplication of the state space by the combinatorial number of internal task configurations leads to an acute curse of dimensionality. Our proposed $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, with dynamic and infinite priority classes, introduces even more complex scheduling dilemmas.
In addition, the multi-task job structure raises a critical question of priority assignment: should priority be an attribute of the entire job (job level priority) or of its individual constituent tasks (task level priority)? A job level policy, where all tasks of a high-priority job inherit its priority, risks resource monopolization if tasks can be parallelized, potentially starving all other jobs. Conversely, a task level policy could cause a critical job to stall if one of its non-essential, low-priority tasks is indefinitely delayed. In this paper, our analysis demonstrates that assigning priority at the job level is, perhaps counter-intuitively, effective in an asymptotic sense.
It is worth mentioning that the seminal work by Leonardi and Raz [23] introduced a general technique to transform any preemptive scheduling solution into a non-preemptive one, incurring an additional approximation factor of $O(\sqrt{n/m})$ for total flow time, where $n$ is the number of jobs and $m$ is the number of machines. The approximation ratio is computed by evaluating the increase in total flow time due to the transformation from a preemptive to a non-preemptive solution. However, this transformation relies on complete knowledge of the job set, i.e., total flow time in the preemptive solution, to do job classification, making it inherently offline and limiting its applicability where jobs arrive dynamically. In contrast, our proposed $\mathsf{NP\text{-}SRPT}$ algorithm extends the transformation of preemptive SRPT into a non-preemptive framework in an online-compatible manner.
I-C Related Work
Single and multiple task scheduling
@@ I-C Related Work
There has been a large literature on single-task job scheduling, with parallel developments taking place in competitive analysis and queuing theory. For example, recently Hong and Scully [15] showed Gittins’s heavy-traffic optimality in $\mathsf{G}/\mathsf{G}/\mathsf{N}$. However, little is known about multitask scheduling. Scully et. al [36] presented the first theoretical analysis of single-processor multitask scheduling problem, and gave an optimal policy that is easy to compute for batch arrival, together with the assumption that the processing time of tasks satisfies the aged Pareto distributions. Sun et al. [43] studied the multitask scheduling problem when all the tasks are of unit size, and proved that among causal and non-preemptive policies, fewest unassigned tasks first (FUT) policy, earliest due date first (EDD) policy, and first come first serve (FCFS) are near delay-optimal in distribution (stochastic ordering) for minimizing the metric of average delay, maximum lateness and maximum delay respectively. Wang et. al [44] established results of asymptotic independence of queues for multitask parallel jobs, by developing a new technique named Poisson oversampling. To model the scenario when the scheduler has incomplete information about the job size, Scully et. al [49] introduced the multistage job model and proposed an optimal scheduling algorithm for multistage job scheduling in $\mathsf{M}$$\mathsf{G}$/11 queue. The closed-form expression of the mean response time is also given for the optimal scheduler. In addition to the aforementioned work, there are also studies that further extend the understanding of scheduling by considering parallelizable jobs represented as Directed Acyclic Graphs (DAGs) [1].
Advances in non-preemptive scheduling
Non-preemptive scheduling, characterized by the irrevocable assignment of jobs to uninterrupted execution slots, encounters profound analytical challenges stemming from state-space explosion and the absence of flexible interruption mechanisms, as evidenced in both adversarial and stochastic settings. These difficulties manifest in stringent lower bounds, such as $\Omega(n^{1/2-\varepsilon})$-inapproximability for minimizing total flow-time even on a single machine  [21] and $\Omega(n^{1/3-\varepsilon})$ on parallel machines [23]. Several algorithmic results [17, 31, 29, 30] circumvent such barriers through relaxed models like resource augmentation, where algorithms access faster processors than adversaries, and rejection, permitting the discard of a small job fraction.
Non-preemptive multiserver priority queues are considered in queueing theory, which has resisted general exact analysis except in highly constrained regimes, and has led to a long line of special-case solutions and approximations. Exact results are obtained only in very restricted cases, such as identical service-time distributions across classes [9] or systems with only two priority levels [47, 22, 50]. Similarly, restrictions to two priority levels have enabled precise computations of performance metrics like waiting times and queue lengths, with recent advancements deriving closed-form expressions via recurrence relations and contour integrals for equal service rates [50]. When service times differ, results like those of Gail et al. [10] suffer from numerical instability as the number of servers grows. Several approaches exploited structured state-space methods, e.g., quasi-birth-and-death formulations[20] or matrix-geometric solutions [35], but these remain feasible only for very small systems. To extend applicability, approximation approaches have emerged for complex arrival patterns, such as quasi-Poisson processes with multiple priority classes [7].
Applications in modern computing further illustrate progress, practical extensions in cloud environments have adapted non-preemptive models for virtual machine scheduling, leveraging knapsack-based algorithms to balance throughput and delay [34]. Investigations into capacity rationing in multiserver nonpreemptive priority queues, where low-priority access is restricted unless idle servers exceed a threshold, provide asymptotic insights for large-scale systems with closely matched supply and demand [6].
All these contributions advance performance analysis, but also underscore the need for methods in non-preemptive setting to handle diverse service distributions and priority structures.
Performance and optimality of SRPT and its variants
@@
While SRPT is optimal at minimizing average response time in single-server systems [41], its performance is suboptimal in multi-server environments. However, SRPT remains a highly regarded method in the context of multi-server systems. It has shown that the competitive ratio of SRPT is in logarithmic order and is the best possible [23]. With the stochastic arrivals in $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ systems, SRPT is shown to be optimal in heavy traffic scenarios [12]. Another notable contribution by [18] introduces the Intermediate-SRPT algorithm, which addresses jobs with intermediate parallelizability and establishes an optimal $O(\log P)$-competitive ratio concerning average flow time, where $P$ represents the ratio of maximum to minimum job sizes.
I-D Paper organization
@@
The remainder of this paper is organized as following. We introduce the problem definition, notations and necessary background in Section II. In Section III we formally present $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, together with the analysis of its competitive ratio and lower bounds. Section IV is devoted to the proof of the asymptotic optimality of $\mathsf{NP}$-$\mathsf{SRPT}$ in heavy traffic regime, and the optimality is extended to infinite task size in Section V. In Section VI, the analysis is extended to the setting of unknown job sizes. We conduct the experimental validation in Section VII, and conclude our work in Section VIII.
II Model and preliminaries
Deterministic Model
@@ II Model and preliminaries
Throughout the paper we use $\alpha=\max_{i\in[n]}{p_{i}}/\min_{i\in[n]}{p_{i}}$ to denote the ratio of the maximum to the minimum job workload. Let $\eta=\max\{p_{i,\ell}|\mbox{ task }\ell\mbox{ of job }i\mbox{ is non-preemptive}\}$ be the maximum processing time of a non-preemptive task, $\beta=\eta/\min_{i\in[n]}{p_{i}}$ be the ratio between $\eta$ and minimum job workload. In some sense, parameters $\beta$ and $\eta$ represent the degree of non-preemptivity and exhibits a trade-off between the preemptive and non-preemptive setting. The problem approaches the preemptive case when $\eta$ is small, and degenerates to the non-preemptive case if all the jobs are consisted of a single non-preemptive task, in which $\eta$ reaches the maximum value of $\max_{i\in[n]}{p_{i}}$.
@@
Stochastic Model
In the stochastic setting, we assume that jobs arrive into the system according to a Poisson process with rate $\lambda$. Job processing times are i.i.d distributed with probability density function $f(\cdot)$. Formally, we consider a sequence of $\mathsf{M}$$\mathsf{G}$$\mathsf{N}$ queues indexed by $k$, the traffic intensity of the $k$-th system is equal to $\rho^{(k)}=\lambda^{(k)}\cdot\mathbbm{E}[p^{(k)}_{i}$], where $\lambda^{(k)}$ denotes the arrival rate of the $k$-th Poisson arrival process, job workload distribution has a density function of $f^{(k)}(\cdot)$. Stability of the queuing systems requires that $\rho^{(k)}<1$ for $\forall k$. As standard in the literature, we assume that $\rho^{(k)}\rightarrow 1$ when $k\rightarrow\infty$. In this paper, we further assume that the probability density function $f^{(k)}(\cdot)$ is continuous. For notational convenience, we will suppress index $n$ whenever it is clear from the context.
The stochastic analysis in this paper relies heavily on the concept of busy period in $\mathsf{M}$$\mathsf{G}$/11, which is defined as following.
Definition 3 (Busy Period in $\mathsf{M}$$\mathsf{G}$/11 [13]).
Busy period in $\mathsf{M}$$\mathsf{G}$/11 is defined to be the longest time interval in which no machines are idle.
We use $\mathsf{B}(w)$ to denote the expected length of a busy period in $\mathsf{M}$$\mathsf{G}$/11, started by a workload of $w$. It can be seen that $\mathsf{B}(\cdot)$ is an additive function [13], i.e., $\mathsf{B}(w_{1}+w_{2})=\mathsf{B}(w_{1})+\mathsf{B}(w_{2})$ for any independent initial workload $w_{1},w_{2}$, since a busy period in $\mathsf{M}$$\mathsf{G}$/11 with initial workload of $w_{1}+w_{2}$ can be regarded as a busy period started by initial workload $w_{2}$, following a busy period started by initial workload $w_{1}$. Moreover, for $\mathsf{M}$$\mathsf{G}$/11 queue, the length of a busy period with initial workload of $w$ and load $\rho$ is shown to be equal to $\mathsf{B}(w)=\mathbbm{E}[w]/(1-\rho)$ [13].
SOAP policy
@@
A SOAP policy is monotonic if its rank function is monotonic in age $a$. In this paper, a tie-breaking rule, First-Come First-Served (FCFS) is applied to determine the order of service in the event of a tie. We use $p$ to denote the specific descriptor $d$ for our size-based ploicy.
III $\mathsf{NP}$-$\mathsf{SRPT}$ Algorithm and Competitive Ratio Analysis